# 3D-LFM: Lifting Foundation Model

### v1

@@ Abstract
The lifting of 3D structure and camera from 2D landmarks is at the cornerstone of the entire discipline of computer vision. Traditional methods have been confined to specific rigid objects, such as those in Perspective-n-Point (PnP) problems, but deep learning has expanded our capability to reconstruct a wide range of object classes (e.g. C3PDO [18] and PAUL [24]) with resilience to noise, occlusions, and perspective distortions. All these techniques, however, have been limited by the fundamental need to establish correspondences across the 3D training data – significantly limiting their utility to applications where one has an abundance of “in-correspondence” 3D data. Our approach harnesses the inherent permutation equivariance of transformers to manage varying number of points per 3D data instance, withstands occlusions, and generalizes to unseen categories. We demonstrate state of the art performance across 2D-3D lifting task benchmarks. Since our approach can be trained across such a broad class of structures we refer to it simply as a 3D Lifting Foundation Model (3D-LFM) -– the first of its kind.
@@ 1 Introduction
Lifting 2D landmarks from a single-view RGB image into 3D has long posed a complex challenge in the field of computer vision because of the ill-posed nature of the problem. This task is important for a range of applications from augmented reality to robotics, and requires an understanding of non-rigid spatial geometry and accurate object descriptions [11, 25, 2]. Historically, efforts in single-frame 2D-3D lifting have encountered significant hurdles: a reliance on object-specific models, poor scalability, and limited adaptability to diverse and complex object categories. Traditional methods, while advancing in specific domains like human body [14, 31, 16] or hand modeling [3, 6], often fail when faced with the complexities of varying object types or object rigs (skeleton placements).
@@
To facilitate such single-frame 2D-3D lifting, deep learning methods like C3DPO [18] and others [11, 24, 8, 28, 25] have been recently developed. However, these methods are fundamentally limited in that they must have knowledge of the object category and how the 2D landmarks correspond semantically to the 2D/3D data it was trained upon. Further, this represents a drawback, especially when considering their scaling up to dozens or even hundreds of object categories, with varying numbers of landmarks and configurations. This paper marks a departure from such correspondence constraints, introducing the 3D Lifting Foundation Model (3D-LFM), an object-agnostic single frame 2D-3D lifting approach. At its core, the 3D-LFM addresses the limitation of previous models, which is the inability to efficiently handle a wide array of object categories while maintaining high fidelity in 3D keypoint lifting from 2D data. We propose a solution rooted in the concept of permutation equivariance, a property that allows our model to autonomously establish correspondences among diverse sets of input 2D keypoints.
@@
3D-LFM is able to carry out single-frame 2D-3D lifting for $30+$ categories using a single model simultaneously, covering everything from human forms [32, 15, 9], face [29], hands [17], and animal species [10, 1, 27], to a plethora of inanimate objects found in everyday scenarios such as cars, furniture, etc. [26]. 3D-LFM is able to achieve 2D-3D lifting performance that matches those of leading methods specifically optimized for individual categories. 3D LFM’s generalizability is further evident in its ability to handle out-of-distribution (OOD) object categories and rigs, which we refer as OOD 2D-3D lifting, where the task is to lift the 2D landmarks to 3D for a category never seen during training. We show such OOD results: (1) for inanimate objects - by holding out an object category within the PASCAL dataset, (2) for animals - by training on common object categories such as dogs and cats found in [27] and reconstructing 3D for unseen and rare species of Cheetahs found in [10] and in-the-wild zoo captures from [5], and (3) by showing rig transfer, i.e. training 2D to 3D lifting on a Human3.6M dataset rig [7] and showing similar 2D to 3D lifting performance on previously unseen rigs such as those found in Panoptic studio dataaset rig [9] or a COCO dataset rig [13]. 3D-LFM transfers learnings from seen data during training to unseen OOD data during inference. It does so by learning general structural features during the training phase via the proposed permutation equivariance properties and specific design choices that we discuss in the following sections.
@@
Recognizing the important role that geometry plays in 3D reconstruction [24, 18, 5, 4, 11, 25], we integrate Procrustean methods such as Orthographic-N-Point (OnP) or Perspective-N-Point (PnP) to direct the model’s focus on deformable aspects within a canonical frame. This incorporation significantly reduces the computational onus on the model, freeing it from learning redundant rigid rotations and focusing its capabilities on capturing the true geometric essence of objects. Scalability, a critical aspect of our model, is addressed through the use of tokenized positional encoding (TPE), which when combined with graph-based transformer architecture, not only enhances the model’s adaptability across diverse categories but also strengthens its ability to handle multiple categories with different number of keypoints and configurations. Moreover, the use of skeleton information (joint connectivity) within the graph-based transformers via adjacency matrices provides strong clues about joint proximity and inherent connectivity, aiding in the handling of correspondences across varied object categories.
@@
    The integration of tokenized positional encoding within the transformer, to enhance our approach’s scalability and its capacity to handle diverse and imbalanced datasets.
@@
    We demonstrate that 3D-LFM surpasses state-of-the-art methods in categories like humans, hands, and faces (benchmark in [32]). Additionally, it shows robust generalization by handling previously unseen objects and configurations, including animals ([10, 5]), inanimate objects ([26]), and novel object arrangements (rig transfer in [9])
In subsequent sections, we explore the design and methodology of our proposed 3D-LFM architecture, including detailed ablation experiments and comparative analyses. Throughout this paper, ’keypoints,’ ’landmarks,’ and ’joints’ are used interchangeably, referring to specific, identifiable points or locations on an object or figure that are crucial for understanding its structure and geometry.
@@ 2 Related works
The field of 2D-3D lifting has evolved substantially from classic works such as the Perspective-n-Point (PnP) algorithm [12]. In these early works, the algorithm was given a set of 2D landmarks and some 3D supervision – namely the known 3D rigid object. The field has since witnessed a paradigm shift with the introduction of deep learning methodologies, led by methods such as C3DPO [18], PAUL [24], and Deep NRSfM [11], along with recent transformer-based innovations like NRSfMFormer [8]. In these approaches one does not need knowledge of the specific 3D object, instead it can get away with just the 2D landmarks and correspondences to an ensemble of 2D/3D data from the object category to be lifted. However, despite their recent success, all these methods still require that the 2D/3D data be in semantic correspondence. That is the index to a specific landmark has the same semantic meaning across all instances (e.g. chair leg). In practice, this is quite limiting at run-time, as one needs intimate knowledge of the object category, and rig in order to apply any of these current methods. Further, this dramatically limits the ability of these methods to leverage cross-object and cross-rig datasets, prohibiting the construction of a truly generalizable 2D to 3D foundation lifting model – a topic of central focus in this paper.
Recent literature in pose estimation, loosely connected to NRSfM but often more specialized towards human and animal body part, has also seen remarkable progress. Models such as Jointformer [14] and SimpleBaseline [16] have refined the single-frame 2D-3D lifting process, while generative approaches like MotionCLIP [19] and Human Motion Diffusion Model [20] have laid the groundwork for 3D generative motion-based foundation models. These approaches, however, are even more limiting than C3PDO, PAUL, etc. in that they are intimately wedded to the object class and are not easily extendable to an arbitrary objects class.
@@ 3 Approach
The 3D lifting function $f:\mathbb{R}^{N\times 2}\rightarrow\mathbb{R}^{N\times 3}$ maps the 2D keypoints to their corresponding 3D structure while compensating for the orthographic projection:
$\mathbf{S}=f(\mathbf{W})=\mathbf{W}\mathbf{P}^{\top}+\mathbf{b}$ (2)
where $\mathbf{P}\in\mathbb{R}^{3\times 2}$ is the orthographic projection matrix and $\mathbf{b}\in\mathbb{R}^{N\times 3}$ is a bias term that aligns the centroids of 2D and 3D keypoints.
@@
where $D$ is the dimensionality of the Fourier feature space, $\boldsymbol{\omega}\in\mathbb{R}^{2\times\frac{D}{2}}$ and $\mathbf{b}\in\mathbb{R}^{\frac{D}{2}}$ are parameters sampled from a normal distribution, scaled appropriately. These parameters are sampled once and kept fixed, as per the RFF methodology. The output of this transformation $\mathbf{TPE}(\mathbf{W}_{c})$ is then fed into the Graph Transformer network as $\mathbf{X}^{\ell}$ where $\ell$ indicates the layer number (00 in the above case). This set of features is now ready for processing inside the graph-based transformer layers without the need for correspondence among the input keypoints. The TPE retains the permutation equivariance property while implicitly encoding the relative positions of the keypoints.
@@ 3.1 Graph-based Transformer Architecture
Hybrid Feature Aggregation: For each layer $\ell$, with $\ell$ ranging from 0 to $L$, the feature matrix $\mathbf{X}^{(\ell)}\in\mathbb{R}^{N\times D}$ is augmented through simultaneous local and global processing:
@@
Here, GA represents Graph Attention, MHSA denotes Multi-Head Self-Attention, and MLP_GeLU indicates our MLP with GeLU nonlinearity. This architecture is designed to learn complex patterns by considering both the local neighborhood connectivity of input 2D and the global data context of input 2D, which is important for robust 2D to 3D structure lifting.
@@ 3.2 Procrustean Alignment
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}=\text{SVD}((\mathbf{M}\odot\mathbf{S}_{c})^{\top}\mathbf{S}_{r})$
$\mathbf{R}=\mathbf{U}\mathbf{V}^{\top}$
@@ 4 Results and Comparative Analysis
Our evaluation shows the 3D Lifting Foundation Model (3D-LFM)’s capability in single-frame 2D-3D lifting across diverse object categories without object-specific data in Sec. 4.1. Following that, Sec. 4.2 highlights 3D-LFM’s performance over specialized methods, especially in whole-body benchmarks (referenced in [32]), showcasing adaptability across varied categories like the human body, face, and hands (Fig.3). Additionally, Section4.3 shows 3D-LFM’s capability in 2D-3D lifting across 30 categories using a single unified model, enhancing category-specific performance and achieving out-of-distribution (OOD) generalization for unseen object configurations during training.
Concluding, ablation studies in Section 4.4 validates our proposed OnP approach, token positional encoding, and the hybrid-attention mechanism in the transformer model, confirming their role in 3D-LFM’s effectiveness in both single and multiple-object scenarios.
@@ 4.1 Multi-Object 3D Reconstruction
Experiment Rationale: The 3D-LFM is designed to process diverse sets of 2D keypoints and lift them into 3D across multiple object categories without relying on category-specific information. Its key strength lies in permutation equivariance, which maintains consistent performance regardless of keypoint order—this is critical when handling multiple objects with different numbers of keypoints. Unlike methods that depend on fixed-dimension arrays to manage object information, our model adapts flexibly to the data at hand. It has been evaluated against non-rigid structure-from-motion approaches [18, 24, 11, 25] that require object-specific inputs, showing its ability to handle diverse categories. For a comprehensive benchmark, we utilize the PASCAL3D+ dataset [26], following C3DPO’s [18] methodology, to include a variety of object categories.
Performance: We benchmark our 3D-LFM against C3DPO [18], a leading method in the NRSfM domain, to evaluate its performance in multi-object 2D to 3D lifting tasks, with 3D supervision. While other recent methods [24, 28, 11, 25] also require object-specific details, C3DPO’s approach to multiple object categories makes it a suitable comparison for our model. Initially, we replicate conditions where object-specific information is provided, resulting in comparable performance between 3D-LFM and C3DPO, evident in Fig. 2. This stage involves using the Mean-per-joint-position-error (MPJPE) to measure 2D-3D lifting accuracy, with C3DPO’s training setup including an $MN$ dimensional array for object details where $M$ represents number of objects with $N$ being maximum number of keypoints, and our model is trained separately on each object to avoid avoid providng object-specific information.
The 3D-LFM’s strength is most apparent when object-specific data is withheld. Unlike C3DPO, which experiences performance drops without object details, our model sustains a lower MPJPE, even when trained collectively across categories using only an $N$ dimensional array. The results (Fig.2 and Tab.1) highlight 3D-LFM’s robustness and superior adaptability, outperforming single-category training and demonstrating its potential as a generalized 2D to 3D lifting solution.
Figure 2: This figure shows the MPJPE for 3D-LFM and C3DPO using the PASCAL3D+ dataset, comparing performance with and without object-specific information. Both methods perform comparably when object-specific information is provided. However, absent this data, C3DPO’s error increases markedly, whereas 3D-LFM’s remains low, showcasing its robustness and adaptability across varied object categories without relying on object-specific information.
Table 1: Quantitative Comparison of Multi-Object 2D-3D Lifting Performance
Method Object-specific MPJPE (avg)
C3DPO [18] Yes $7.5$
3D-LFM (Ours) Yes $\mathbf{3.97}$
C3DPO [18] No $41.08$
3D-LFM (Ours) No $\mathbf{3.27}$

### v2

@@ Abstract
The lifting of a 3D structure and camera from 2D landmarks is at the cornerstone of the discipline of computer vision. Traditional methods have been confined to specific rigid objects, such as those in Perspective-n-Point (PnP) problems, but deep learning has expanded our capability to reconstruct a wide range of object classes (e.g. C3DPO [18] and PAUL [24]) with resilience to noise, occlusions, and perspective distortions. However, all these techniques have been limited by the fundamental need to establish correspondences across the 3D training data, significantly limiting their utility to applications where one has an abundance of “in-correspondence” 3D data. Our approach harnesses the inherent permutation equivariance of transformers to manage varying numbers of points per 3D data instance, withstands occlusions, and generalizes to unseen categories. We demonstrate state-of-the-art performance across 2D-3D lifting task benchmarks. Since our approach can be trained across such a broad class of structures, we refer to it simply as a 3D Lifting Foundation Model (3D-LFM) – the first of its kind.
@@ 1 Introduction
Lifting 2D landmarks from a single-view RGB image into 3D has long posed a complex challenge in the field of computer vision because of the ill-posed nature of the problem. This task is important for a range of applications from augmented reality to robotics, and requires an understanding of non-rigid spatial geometry and accurate object descriptions [11, 25, 2]. Historically, efforts in single-frame 2D-3D lifting have encountered significant hurdles: reliance on object-specific models, poor scalability, and limited adaptability to diverse and complex object categories. Traditional methods, while advancing in specific domains like human body [14, 31, 16] or hand modeling [3, 6], often fail when faced with the complexities of varying object types or object rigs (skeleton placements).
@@
To facilitate such single-frame 2D-3D lifting, deep learning methods like C3DPO [18] and others [11, 24, 8, 28, 25] have recently been developed. However, these methods are fundamentally limited in that they must have knowledge of the object category and how the 2D landmarks correspond semantically to the 2D/3D data it was trained upon. Further, this represents a drawback, especially when considering their scaling up to dozens or even hundreds of object categories, with varying numbers of landmarks and configurations. This paper marks a departure from such correspondence constraints, introducing the 3D Lifting Foundation Model (3D-LFM), an object-agnostic single frame 2D-3D lifting approach. At its core, 3D-LFM addresses the limitation of previous models, which is the inability to efficiently handle a wide array of object categories while maintaining high fidelity in 3D keypoint lifting from 2D data. We propose a solution rooted in the concept of permutation equivariance, a property that allows our model to autonomously establish correspondences among diverse sets of input 2D keypoints.
@@
3D-LFM is capable of performing single frame 2D-3D lifting for $30+$ categories using a single model simultaneously, covering everything from human forms [32, 15, 9], face [29], hands [17], and animal species [10, 1, 27], to a plethora of inanimate objects found in everyday scenarios such as cars, furniture, etc. [26]. Importantly, 3D-LFM is inherently scalable, poised to expand to hundreds of categories and improve performance, especially in out-of-distribution or less-represented areas, showcasing its broad utility in 3D lifting tasks. 3D-LFM is able to achieve 2D-3D lifting performance that matches those of leading methods specifically optimized for individual categories. The generalizability of 3D LFM is further evident in its ability to handle out-of-distribution (OOD) object categories and rigs, which we refer to as OOD 2D-3D lifting, where the task is to lift the 2D landmarks to 3D for a category never seen during training. We show such OOD results: (1) for inanimate objects - by holding out an object category within the PASCAL dataset, (2) for animals - by training on common object categories such as dogs and cats found in [27] and reconstructing 3D for unseen and rare species of Cheetahs found in [10] and in-the-wild zoo captures from [5], and (3) by showing rig transfer, i.e. training 2D to 3D lifting on a Human3.6M dataset rig [7] and showing similar 2D to 3D lifting performance on previously unseen rigs such as those found in Panoptic studio dataaset rig [9] or a COCO dataset rig [13]. 3D-LFM transfers learnings from seen data during training to unseen OOD data during inference. It does so by learning general structural features during the training phase via the proposed permutation equivariance properties and specific design choices that we discuss in the following sections.
@@
Recognizing the important role geometry plays in 3D reconstruction [24, 18, 5, 4, 11, 25], we integrate Procrustean methods such as Orthographic-N-Point (OnP) or Perspective-N-Point (PnP) to direct the model’s focus on deformable aspects within a canonical frame. This incorporation significantly reduces the computational burden on the model, freeing it from learning redundant rigid rotations and focusing its capabilities on capturing the true geometric essence of objects. Scalability, a critical aspect of our model, is addressed through the use of tokenized positional encoding (TPE), which, when combined with graph-based transformer architecture, not only enhances the model’s adaptability across diverse categories but also strengthens its ability to handle multiple categories with different number of keypoints and configurations. Finally, the use of skeleton information (joint connectivity) within the graph-based transformers via adjacency matrices provides strong clues about joint proximity and inherent connectivity, aiding in the handling of correspondences across varied object categories.
@@
    The integration of tokenized positional encoding within the graph-based transformer, to enhance our approach’s scalability and its capacity to handle diverse and imbalanced datasets.
@@
    We demonstrate that 3D-LFM surpasses state-of-the-art methods in categories such as humans, hands, and faces (benchmark in [32]). Additionally, it shows robust generalization by handling previously unseen objects and configurations, including animals ([10, 5]), inanimate objects ([26]), and novel object arrangements (rig transfer in [9])
In subsequent sections, we explore the design and methodology of our proposed 3D-LFM architecture, including detailed ablation experiments and comparative analyses. Throughout this paper, ‘keypoints’, ‘landmarks’, and ‘joints’ are used interchangeably, referring to specific, identifiable points or locations on an object or figure that are crucial for understanding its structure and geometry.
@@ 2 Related works
The field of 2D-3D lifting has evolved substantially from classic works such as those based on Perspective-n-Point (PnP) algorithms [12]. In these early works, the algorithm was given a set of 2D landmarks and some 3D supervision, namely the known 3D rigid object. The field has since witnessed a paradigm shift with the introduction of deep learning methodologies, led by methods such as C3DPO [18], PAUL [24], and Deep NRSfM [11], along with recent transformer-based innovations such as NRSfMFormer [8]. In these approaches one does not need knowledge of the specific 3D object, instead it can get away with just the 2D landmarks and correspondences to an ensemble of 2D/3D data from the object category to be lifted. However, despite their recent success, all these methods still require that the 2D/3D data be in semantic correspondence. That is, the index to a specific landmark has the same semantic meaning across all instances (e.g. chair leg). In practice, this is quite limiting at run-time, as one needs intimate knowledge of the object category, and rig in order to apply any of these current methods. Further, this dramatically limits the ability of these methods to leverage cross-object and cross-rig datasets, prohibiting the construction of a truly generalizable 2D to 3D lifting foundation model – a topic of central focus in this paper.
Recent literature in pose estimation, loosely connected to NRSfM but often more specialized towards human and animal body parts, has also seen remarkable progress. Models such as Jointformer [14] and SimpleBaseline [16] have refined the single-frame 2D-3D lifting process, while generative approaches like MotionCLIP [19] and Human Motion Diffusion Model [20] have laid the groundwork for 3D generative motion-based foundation models. These approaches, however, are even more limiting than C3PDO, PAUL, etc. in that they are intimately wedded to the object class and are not easily extendable to an arbitrary object class.
@@ 3 Approach
The 3D lifting function $f:\mathbb{R}^{N\times 2}\rightarrow\mathbb{R}^{N\times 3}$ maps the 2D keypoints to their corresponding 3D structure while compensating for the projection:
$\mathbf{S}=f(\mathbf{W})=\mathbf{W}\mathbf{R}^{\top}+\mathbf{b}$ (2)
where $\mathbf{R}\in\mathbb{R}^{3\times 3}$ is the projection matrix (assumed either weak-perspective or orthographic) and $\mathbf{b}\in\mathbb{R}^{N\times 3}$ is a bias term that aligns the centroids of 2D and 3D keypoints.
@@
where $D$ is the dimensionality of the Fourier feature space, $\boldsymbol{\omega}\in\mathbb{R}^{2\times\frac{D}{2}}$ and $\mathbf{b}\in\mathbb{R}^{\frac{D}{2}}$ are parameters sampled from a normal distribution, scaled appropriately. These parameters are sampled once and kept fixed, as per the RFF methodology. The output of this transformation $\mathbf{TPE}(\mathbf{W}_{c})$ is then fed into the graph-based transformer network as $\mathbf{X}^{\ell}$ where $\ell$ indicates the layer number (00 in the above case). This set of features is now ready for processing inside the graph-based transformer layers without the need for correspondence among the input keypoints. The TPE retains the property of permutation equivariance while implicitly encoding the relative positions of the keypoints.
@@ 3.1 Graph-based Transformer Architecture
Hybrid Feature Aggregation: For each layer $\ell$, ranging from 0 to $L$, the feature matrix $\mathbf{X}^{(\ell)}\in\mathbb{R}^{N\times D}$ is augmented through simultaneous local and global processing. The local processing component, GA($\mathbf{X}^{(\ell)},\mathbf{A}$), leverages an adjacency matrix $\mathbf{A}$, which encodes the connectivity based on the object category, to perform graph-based attention on batches of nodes representing the input 2D data:
@@
Here, GA represents Graph Attention, MHSA denotes Multi-Head Self-Attention, and MLP_GeLU indicates our MLP with GeLU nonlinearity. This architecture is designed to learn patterns in 2D data by considering both the local neighborhood connectivity of input 2D and the global data context of input 2D, which is important for robust 2D to 3D structure lifting.
@@ 3.2 Procrustean Alignment
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}=\text{SVD}((\mathbf{M}\odot\mathbf{S}_{c})^{\top}\mathbf{S}_{r}),\quad\mathbf{R}=\mathbf{U}\mathbf{V}^{\top}$
@@ 4 Results and Comparative Analysis
Our evaluation shows the 3D Lifting Foundation Model (3D-LFM)’s capability in single-frame 2D-3D lifting across diverse object categories without object-specific data in Sec. 4.1. Following that, Sec. 4.2 highlights 3D-LFM’s performance over specialized methods, especially achieving state-of-the-art performance in whole-body benchmarks[32] (Fig. 3). Additionally, Sec. 4.3 shows 3D-LFM’s capability in 2D-3D lifting across 30 categories using a single unified model, enhancing category-specific performance and achieving out-of-distribution (OOD) generalization for unseen object configurations during training. In conclusion, the ablation studies in Section 4.4 validate our proposed procrustean approach, token positional encoding, and the local-global hybrid attention mechanism in the transformer model, confirming their role in 3D-LFM’s effectiveness in both single- and multiple-object scenarios.
@@ 4.1 Multi-Object 3D Reconstruction
Clarifying naming convention: In ‘object-specific’ versus ‘object-agnostic’, our primary focus in this naming is on the distinction in training methods. Here, object-specific training involves supplying semantic details for each object, leading to isolated training. Conversely, object-agnostic training combines various categories without explicit landmark semantics, leading to combined training.
Experiment Rationale: 3D-LFM leverages permutation equivariance to accurately lift 2D keypoints into 3D structures across diverse categories, outperforming fixed-array methods by adapting flexibly to variable keypoint configurations. It has been evaluated against non-rigid structure-from-motion approaches [18, 24, 11, 25] that require object-specific inputs, showing its ability to handle diverse categories. For a comprehensive benchmark, we utilize the PASCAL3D+ dataset [26], following C3DPO’s [18] methodology, to include a variety of object categories.
Performance: We benchmark 3D-LFM against the notable NRSfM method, C3DPO [18], for multi-object 2D to 3D lifting with 3D supervision. C3DPO, similar to other contemporary methods [24, 28, 11, 25] requiring object-specific details, serves as an apt comparison due to its multi-category approach. Initially replicating conditions with object-specific information, 3D-LFM matches C3DPO’s performance, as demonstrated in Fig. 2. This stage uses MPJPE to measure 3D lifting accuracy, with C3DPO’s training setup including an $MN$ dimensional array for object details where $M$ represents number of objects with $N$ being maximum number of keypoints, and our model is trained separately on each object to avoid providing object-specific information. The 3D-LFM’s strength emerges when object-specific data is withheld. While C3DPO shows a decline without such data, 3D-LFM maintains a lower MPJPE across categories, even when trained collectively across categories using only an $N$ dimensional array. These findings (Fig. 2) highlights 3D-LFM’s capabilities, outperforming single-category training and demonstrating its potential as a generalized 2D to 3D lifting solution.
Figure 2: 3D-LFM vs. C3DPO Performance: MPJPE comparisons using the PASCAL3D+ dataset, this figure demonstrates our model’s adaptability in the absence of object-specific information, contrasting with C3DPO’s increased error under the same conditions. The analysis confirms 3D-LFM’s superiority across diverse object categories, reinforcing its potential for generalized 2D to 3D lifting.